OpenStudy (anonymous):

Simplify. -12-4.8 x (-2+-8)

OpenStudy (igreen):

\(-12 - 4.8 \cdot (-2 + -8)\) Is this it?

OpenStudy (anonymous):

yep

OpenStudy (igreen):

\(-12 - 4.8 \cdot (-2 + -8)\) Simplify: \(-12 - 4.8 \cdot (-2 - 8)\) Subtract: \(-12 - 4.8 \cdot (-10)\) Multiply: \(-12 + 48\) Add: \(36\)

OpenStudy (anonymous):

thanks!